3. Find a polynomial, P(x), of degree 3 with zeros of 4,1 and −1, if P(0) = 8.


1
Expert's answer
2022-05-11T12:43:59-0400
"P(x)=a(x-4)(x-1)(x+1)"

If "P(0)=8"


"P(0)=a(0-4)(0-1)(0+1)=8=>a=2"

"P(x)=2(x-4)(x-1)(x+1)"

"=(2x-8)(x^2-1)"

"=2x^3-8x^2-2x+8"